A truck takes 1/2 hour more than a car to travels distance of 125 km. If the speed of the truck is 50 km/h, what is the speed of the truck?
I've tried using the speed for formula but I got weird answers like 6 and 2.5

Found 2 solutions by alicealc, Alan3354:
Answer by alicealc(293)

Did you mean the speed of the car?

distance = speed * time

for the truck:
125 = 50 * t
t = 125/50 = 2.5 hours

for the car, because it takes 1/2 hours (= 0.5 hours) less than the truck, it needs 2 hours (2.5 - 0.5 = 2) to travel distance of 125 km.
So, the speed of the car:
125 = speed * 2
speed = 125/2 = 62.5 km/h


Answer by Alan3354(69443)

If the speed of the truck is 50 km/h, what is the speed of the truck?
------------------------
50 km/hr
================
If you want the speed of the car:
The time for the truck is
t = d/r = 125/50 = 2.5 hours
--------------
Time for the car is 1/2 hour less
= 2 hours
Speed of the car is
r = d/t = 125/2
= 62.5 km/hr